Fermi National Laboratory


Accelerator Update for June 10, 2005 - June 13, 2005

Friday June 10
The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Tevatron. (TeV) shot setup on hold due to Main Injector (MI) BPM work, and with MiniBooNE taking beam.

Operations reported at 8:16 AM that technicians had the MI BPMs working.

Operations established combination store 4201 at 10:23 AM with an initial luminosity of 85.7E30. Pbar resumed stacking and NuMI resumed taking beam.

Operations reported at 3:35 PM that beam had been sent to Switchyard (SY) so an expert could conduct an octupole study.

NuMI Horn experts began conducting an hour of Horn current tests at 4:42 PM.

SY studies ended at 5:17 PM.

Operations put the Recycler into shot setup to stash antiprotons at 5:41 PM. The shots began at 6:23 PM.

At 6:35 PM, a Debuncher quadrupole power supply (D:QF) tripped off and wouldn't reset. Operators called in an expert.

The Recycler stash ended at 6:56 PM. NuMI resumed taking beam.

Operations reported at 8:46 PM that D:QF had been repaired.

Pbar resumed stacking at 9:01 PM.

At 9:55 PM, the Booster chopper tripped off and wouldn't reset. Operations called in experts to help with the investigations. This shut off all downstream beam.

At 10 PM, Operations reported that a problem with Booster vacuum valves had gotten worse and that they had called in an expert.

Operations reported at 11:07 PM that beam had returned to Pbar, NuMI, and MiniBooNE. Booster experts reported that they had replaced a power supply for the chopper.

The Booster vacuum expert reported that he had replaced a controls card at 11:52 PM.

Saturday June 11
The midnight shift began with Operations monitoring store 4201 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 12:40 AM, a Debuncher RF station (DRF1-6) alarmed and wouldn't reset. Operators replaced a driver chassis at 1:36 AM, but the spare they had used wasn't working either. It was bypassed and repaired later in the morning.

Operators prepared SY enclosures for access and beamline work at 6:50 AM.

Operations terminated store 4201 at 11:01 AM and put the TeV into shot setup at 11:23 AM.

At 12:38 PM, the high voltage for the I- Source tripped off and wouldn't reset. Operators swapped out the regulator and control modules and had it operating normally by 1:09 PM.

Operations established combination store 4202 at 2:11 PM with an initial luminosity of 99.96E30. Pbar resumed stacking and NuMI resumed taking beam.

At 3:39 PM, a vacuum pump in the P1 (located between MI and the TeV) tripped off. Operators closed the vacuum valves, to let the line recover, but by 4:02 PM, the vacuum problem cascaded to the TeV sector F0 Lambertson. Eventually everything recovered.

Operations began a Recycler stash at 7:06 PM and completed it around 8 PM.

Operations began a Recycler stash at 11:29 PM.

Sunday June 12
The midnight shift began with Operations monitoring store 4202 and the stash, with the Recycler stashing, and with MiniBooNE taking beam.

The recycler stash ended at 12:10 AM. Pbar resumed stacking and NuMI resumed taking beam.

At 2:08 AM, cryo system technicians at CHL reported a cold compressor for TeV sector A0 had tripped off.

At 7:53 AM, Operations reported that a MI vacuum pump (IP606D) tripped off. It would not reset and was left off.

At 11:42 AM, Operations reported that CHL had no power. Operators investigated and discovered that feeders 38 and 47 had tripped off, with 47 tripping off due to a switchgear failure. Operators informed the experiments and then terminated store 4202 at 12 PM. Operators began informing everyone.

High Voltage experts began redistributing power at 2:09 PM to allow technicians access to the feeder 47 switching gear.

Operations put the Recycler into shot setup at 3:02 PM.

Operations reported at 5:49 PM that diesel generators were being installed at CHL. They also reported that the TeV was beginning to warm up.

The Recycler stash ended at 6:37 PM. Pbar resumed stacking and NuMI resumed taking beam.

A High Voltage expert reported at 8:03 PM that they had reenergized feeder 38.

At 8:29 PM, Operations reported an alarm for a helium cold compressor for TeV sector E0. The Cryo Coordinator responded and reported at 9:55 PM that he had turn off the compressor.

MI experts began using antiprotons at 9:09 PM as a tune up for sending particles to the TeV (a fake shot). At 10:12 PM, Operations reported a problem with the fake shots.

A vacuum expert reported at 10:17 PM that all vacuum valves in the TeV had been closed to try and control the vacuum problems caused by the TeV cryo system warm up.

At 10:48 PM, Operation reported that a NuMI C magnet in the MI line had tripped off and wouldn't reset. An operator called in experts.

Operations reported at 11:55 PM that they had prepared the Tevatron for access.

Monday June 13
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with NuMI off due to a ground fault and the TeV off due to a CHL power failure, and with MiniBooNE taking beam.

The NuMI power supply experts reported at 1:16 AM that there was a ground fault on E:V100 in the MI tunnel. (Later, during the access, it was discover that the magnet was okay, but a cable had shorted out.)

Vacuum techs accessed the TeV at 2:41 AM.

At 3:20 AM, Operations reported that the helium system for the Debuncher could not be maintained, so stacking would have to stop, but the stack could be saved.

At 3:22 AM, Operations reported that MI would have to be turned off for an access.

At 3:57 AM, Operations terminated the Recycler stash. Unfortunately, this was the second highest stash that the Recycler had ever held. Operators prepared MI for a tunnel access.

The Plan for Monday
The plans for today to recover from the CHL power outage. TeV personnel will conduct vacuum work and, along with all the other machines, maintenance work will begin. The TeV start up isn't expected to begin before Wednesday evening (June 15), but the other machines may be up by late Monday night (June 13).


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/13/2005   email Fermilab